OpenStudy (anonymous):

HELP PLEASE?! ANYONE ): What term do you need to add to each side of 3x^2 + 15x = 3 to complete the square?

OpenStudy (amistre64):

if we factor out the 3 on the left; we can do this the usual way after wards

OpenStudy (anonymous):

the choices are 25/4 , 15/2, 2/15, or 4/25 ... i had thought the same same thing but...?

OpenStudy (amistre64):

3x^2 + 15x = 3 3(x^2 + 5x) = 3 3((x+n)(x + 5)) = 3 i just factor it as tho its complete, then split the difference between the two parts; n=5/2 3((x+5/2)(x + 5/2)) = 3 3 (x+5/2)^2 = 3 this tells me that the value to complete the square is 3(5/2)^2

OpenStudy (anonymous):

where did u get n from?

OpenStudy (amistre64):

its one of those letter in the alphabet; its just a place filler as i finish it up

OpenStudy (anonymous):

oh okay thnxs for explaining(:

OpenStudy (amistre64):

i get 75/4 which aint a part of your choices tho

OpenStudy (amistre64):

25/4 is which is close, but i cant gaurentee thats the solution tho

OpenStudy (anonymous):

thts alrite i anaylsis it a little more c what i can fine... thanks!

OpenStudy (amistre64):

good luck :)